OpenStudy (anonymous):

Using the number 35, rewrite this number as a difference of two numbers. Show how to use the identity (x − y)2 = x2 − 2xy + y2 to square your number without using a calculator.

OpenStudy (anonymous):

@ranga can you help

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@ganeshie8 Please help

OpenStudy (anonymous):

@iPwnBunnies come on please help

OpenStudy (anonymous):

@ikram002p can you help?

OpenStudy (ipwnbunnies):

Well, what's two numbers that subtract to give you 35?

OpenStudy (anonymous):

70

OpenStudy (ipwnbunnies):

...Wut. If I had the number 10. 13 and 3 would be two numbers that subtract to equal 10.

OpenStudy (ikram002p):

well u just need to right 35 as a diffrence btw two numbers 35=40-5 35=50-15 35=45-10 ...ect

OpenStudy (ikram002p):

to write*

OpenStudy (anonymous):

sorry

OpenStudy (ikram002p):

the second one is asking abt 35^2 if u chose 35=40-5 then it would easy to solve (35)^2=(40-5)^2 can u get it from nw ?

OpenStudy (anonymous):

Um like is that the answer? I have to get this done like in 2 mintutes or i get a F

OpenStudy (ikram002p):

rewrite this number as a difference of two numbers. 35=40-5 Show how to use the identity (x − y)2 = x2 − 2xy + y2 to square your number without using a calculator. (35)^2=(40-5)^2=(40)^2-(5*2*40)+5^2 continue...

OpenStudy (anonymous):

So from on there that is right?

OpenStudy (ikram002p):

its right but u need to continue to find 30^2 (35)^2=(40-5)^2=(40)^2-(5*2*40)+5^2 =1600-400+25 =?